OpenStudy (anonymous):

(Help?) Find the equation of the parabola with focus (2, 3) and directrix y = -1. A) (y - 1)^2 = 8(x - 2) B) (y - 1)^2 = 4(x - 2) C) (x - 2)^2 = 8(y - 1) D) (x - 2)^2 = 4(y - 1)

OpenStudy (anonymous):

\[ (x-2)^2=8 (y-1) \] I will tell you why below

OpenStudy (anonymous):

please do explain.

OpenStudy (anonymous):

The parabola is the set of points equidistant from the focus and directirx

OpenStudy (anonymous):

\[ (y+1)^2=(x-2)^2+(y-3)^2 \]

OpenStudy (anonymous):

Expand and put in the right form.

OpenStudy (anonymous):

\[ y^2+2 y+1=x^2-4 x+y^2-6 y+13 \]

OpenStudy (anonymous):

\[ x^2-4 x+y^2-\left(y^2+2 y+1\right)-6 y+13=x^2-4 x-8 y+12=0 \]

OpenStudy (anonymous):

did you distribute.

OpenStudy (anonymous):

\[ x^2 - 4 x - 8 y + 12 = x^2 - 4 x + 4 + 8 - 8 y = (x - 2)^2 + 8 - 8 y = 0\\ (x - 2)^2 = 8 y-8=8(y-1) \]

OpenStudy (anonymous):

ah never mind i see what you did. That makes tons of sense now to me. Thanks a lot! :)

OpenStudy (anonymous):

YW